My homework asks me to find two positive, consecutive integers whose sum is 624. I think the answer is none, but am not sure that is correct
Answer by nyc_function(2741)

Let x = first integer

Let x + 1 = second integer

x + x + 1 = 624

2x + 1 = 624

2x = 624 - 1

2x = 623

x = 623/2

x = 311.5

The first number is 311.5 and the other is 312.5 and that's it.